OpenStudy (anonymous):

solve for x in the proportion 2x/16x+30=-1/x

hero (hero):

Is it: \(\huge\frac{2x}{16x + 30} = -\frac{1}{x}\)

OpenStudy (anonymous):

yes

hero (hero):

Hint: Cross Multiply